OpenStudy (anonymous):

how many solutions does the system of equations have 2x = -10y + 6 and x + 5y = 3 Can someone show me how to figure this out?

OpenStudy (anonymous):

-4x=-10y

OpenStudy (zenmo):

one second

OpenStudy (anonymous):

alright so wouldn't it be 2x + 1.6=y? or would it stay a negative sign?

OpenStudy (zenmo):

Do you have the solution with you? Could you message me it, I need to check if I'm doing it correct

OpenStudy (anonymous):

i already know the solution i just want to know how to get to the solution (which is infinite )

OpenStudy (zenmo):

okay

OpenStudy (zenmo):

Ok, for the problem, when finding solutions for a system of equations. You want each equation to be in standard form like 2x + 8y = 2

OpenStudy (zenmo):

One of your equations are already in standard form " x + 5y = 3"

OpenStudy (zenmo):

Now, we need 2x = - 10y + 6 to be in standard form, so we need to add +10y to both sides. Now, you get 2x + 10 y = 6

OpenStudy (zenmo):

Are you familiar with elimination?

OpenStudy (anonymous):

no i'm not but i'm trying to learn it

OpenStudy (zenmo):